OpenStudy (anonymous):

The series 2 + 4 + 8 + 16 + 32 + . . .

OpenStudy (anonymous):

If you think of what equation gets you those numbers Think of X^n=# X is the unknown number you are looking for, and n is your position So... Using your series up top X^1=2 X^2=4 X^3=8 If you solve for X you will see it gives you your overall equation.